It is assumed that the number of claims arriving at an insurance company per working day has a mean of 40 and a standard deviation of 12. A survey was conducted over 50 working days. Find the probability that the sample mean number of claims arriving per working day was less than 35.


1
Expert's answer
2022-06-02T12:39:44-0400
"P(X<35)=P(Z<\dfrac{35-40}{12/\sqrt{50}})"

"\approx P(Z<-2.9463)\approx0.0016"